OpenStudy (anonymous):

lim x->16, x^2-256/(sqrt(x)-4)

OpenStudy (anonymous):

first factorize the top into (x+16)(x-16) (difference of squares) then factorize (x-16) into (sqrt(x)-4)(sqrt(x)+4) now cancel and substitute x=16 Answer is 256

OpenStudy (anonymous):

\[{(x^2-256)(\sqrt{x}+4) \over (\sqrt{x}-4)(\sqrt{x}+4)}={(x-16)(x+16)(\sqrt{x}+4) \over x-16}=(x+16)(\sqrt{x}+4)\] Just plug x=16 and get your answer.

OpenStudy (anonymous):

ya thats right ty i did not factor the top first

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

how did you know to factor the top first?

OpenStudy (anonymous):

Because I knew that it would be cancelled out with the bottom.

OpenStudy (anonymous):

haha ok lol i guess i should just do some more

OpenStudy (anonymous):

Yeah I guess so :D